OpenStudy (anonymous):

I have this situation (2+1/x) / (1- 2/x). the answer is 2. How I get there?

OpenStudy (anonymous):

\[ \Large\frac{2+{1\over x}}{1-{2\over x}}=\frac{2x+1\over \cancel{x}}{x-2\over \cancel{x}}\\ \quad={2x+1\over x-2}\color{red}{\ne\Large 2} \]

OpenStudy (anonymous):

I understood. thank you.